OpenStudy (anonymous):

Please Help!?!! :D Which graph represents the solution to the giving system? y=-x+2 and y=3x-1 ...Going to add pictures!

OpenStudy (anonymous):

Add the pictures Amanda Bynes

OpenStudy (anonymous):

Did you add the pictures yet?

OpenStudy (anonymous):

Are you done putting the pictures?

OpenStudy (anonymous):

Hey amanda where are the pictureS?

OpenStudy (anonymous):

Nope my computer frose.

OpenStudy (anonymous):

Froze*

OpenStudy (anonymous):

A B C or D?

OpenStudy (anonymous):

Have you tried looking for the equations on the graphs?

OpenStudy (anonymous):

NOpe??

OpenStudy (anonymous):

y=-x+2 Which graphs have this?

OpenStudy (anonymous):

None

OpenStudy (anonymous):

Okay. -x that tells you the graph wil move right one and down one every time. the +2 means it cross the y intercept at 2 (0,2) Which graphs have this?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

No C has two lines that cross at (0, -2)

OpenStudy (anonymous):

sorry they don't both cross at (0,-2) one does and the other at (0,-1). So try again

OpenStudy (anonymous):

D crosses at (0,-1) that is the closes one I see.

OpenStudy (anonymous):

Yes. Now find the one/s that have a line crossing at (0,2)

OpenStudy (anonymous):

None of them do. I think A is the closes one.

OpenStudy (anonymous):

Okay. Give me the coordinates for the y intercept of each line

OpenStudy (anonymous):

In A.

OpenStudy (anonymous):

1,1

OpenStudy (anonymous):

1,1 is not an intercept. Are you looking at where they cross? It tells you right there (0.75, 1.25) An intercept is when the line cross the x or y axis. So can you give me the y intercepts for both of the lines in A. Hint. (0,x)

OpenStudy (anonymous):

Thanks for your help. :D

OpenStudy (anonymous):

So you know what it is now?

OpenStudy (anonymous):

No. My Brain Just feels like it is going to blow up. I'm just going to do it tomorrow. I have done school work for 7 hours already today.

OpenStudy (anonymous):

This is easy though. I promise. You can do it. Let's just try A.

OpenStudy (anonymous):

I know it is easy I used to be able to do it tell Christmas break. I forgot everything.

OpenStudy (anonymous):

Not good. Okay later babe.

OpenStudy (anonymous):

(0,2)??

OpenStudy (anonymous):

Yes. That's one y-int

OpenStudy (anonymous):

B is (0,2) or (0,-1)

OpenStudy (anonymous):

Yes B Has those y intercepts. Now can you give me the other one for A

OpenStudy (anonymous):

(0,-1)

OpenStudy (anonymous):

Okay. So that's what I said we needed to look for. So now we know that the answer is either A or B. Now let's compare slopes y=-x+2 and y=3x-1 for the line with the y-int of (0,2) the slope is down one for every time you move right one. y=3x-1 for this line the slope has a y int at (0,-1) and it goes up three for ever time you move right one. Which has these characteristics. A or B?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Are you sure about that. To me it looks like the line with a y-int at (0,2) goes up one every time you move right one. Supposed to go down one every time you move right one. and for the line with y-int at (0,-1) I see it goes down three every time you move right one. It's supposed to go up three every time you move right one.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Yeah, so I guess now you know It's A. You did it babe.

OpenStudy (anonymous):

Thank you!